OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price

OpenStudy (anonymous):

Your goal is to find the September price, yes?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Well, take in consideration of the $1.50 afterwards. Since, 20% is an easy percentage to convert and do within your head. So, I took a random amount which was correct. $1.25 20% = 1/5 125/5 = 25. 125 + 25 = 150 (cents) which converts to 1.25 + .25 = 1.50

OpenStudy (anonymous):

If any questions, please ask.

OpenStudy (anonymous):

No I understand Thankyou.

OpenStudy (anonymous):

Of course, anytime. You're welcome. (: